# Can LLMs Grasp Implicit Cultural Values? Benchmarking LLMs’ Metacognitive Cultural Intelligence with CQ-Bench

### v1

@@ Appendix C Prompts
In this section, we show all the prompts we use for story generation, experiments, and validation. The story generation prompt and all post-checking prompts are in Figure C
@@
Story generation prompt You will be provided with 5 cultural values and a location where the conversation happens. Each of them follows the format [culture]–[value]. The first [culture] describes a statement or a situation, and [value] is how you agree with the culture or is the culture common or not. Your task is to generate a scene including conversations and actions among multiple people and the scene needs to reflect the culture values provided. Here are some requirements of the scene: 1. It cannot be too short. It should have multiple rounds of interaction among people. 2. It should not be too obvious. It cannot directly spit out or rephrase the values. 3. It cannot be easy to human to understand the culture values behind. 4. You do not need to follow the order of the values. You could mention the values multiple times through the conversation. Make sure the conversation flows well. 5. All the characters should follow the given values. There should not be contradictions between the character’s value and the given value. Here are the cultural values you should follow when generating: values Here is the pre-defined location of the scene: location Now using the cultural values to generate a story.
[…]
figureStory generation prompt
@@
Incorporation check prompt You will be provided a story and values reflected in the story. Your task is to check if the story reflects values? The story does not mention the values directly. You will need some reasoning to analyze the story. Here is the original story: story Here are the values to reflect in the story: values For each value, output if the value is reflected and provide reasoning. In the end, output the values not reflected without the reasoning. Only output the exact and comprehensive value including ”–” within it, do not rephrase! If all the values are reflected, just leave it blank. Follow the format: [Value]:[Reasoning, Yes/No] ….. Values not reflected: [Value]
[…]
figureIncorporation check prompt
@@
Missing values incorporation prompt You will be provided a story and values which need to be reflected in the story. Your task is to refine the story to reflect the value provided. You cannot remove anything or replace existing speeches from the story, you can only add conversations to reflect the value. The refinement should flow with original story well. You cannot add new conversation randomly. Here is the original story: story Here are the values to reflect in the story: values Now refine the story.
[…]
figureMissing values incorporation prompt
@@
Consistency check prompt You will be provided a story and values reflected in the story. Your task: for each value, check if all the characters agree with the value. If there is characters who does not agree with the value, you should output the character’s name and his speech, and why the speech does not align with the value. Here is the original story: story Here are the values to reflect in the story: values Now check the story and output if there is any contradiction. You can output reasoning to help you analyze. However, in the end, only output where is the contradiction one by one. If there isn’t a contradiction, just reply NO. Otherwise, reply where is the contradiction. Follow the format strictly, do not change the format, output exact values from the values provided, and do not rephrase: [Reasoning]: [Value–attitude]:[If all the speeches are aligned with the value] [Contradictions]: [Value–attitude]: [*character name*:speech] ……
[…]
figureConsistency check prompt
@@
Consistency resolve prompt You will be provided a story and values which need to be reflected in the story. However, the story includes some contradiction where characters do not agree on certain values. You will be provided where is the contradiction. The contradiction includes 3 parts: 1. Correct value to follow 2. Character name 3. Character speech Your task is to replace the speech mentioned in the contradictions with a new speech to make sure the speech is aligned with the values The refinement should flow with the original story well. You cannot add new conversations randomly. Here is the original story: story Here are the contradictions: contradiction Ignore the original character’s speech. Directly write a new speech that reflects the value. Here is the rewritten speech:
[…]
figureConsistency resolve prompt
@@
Implicitness check prompt You will be provided a story and values reflected in the story. Your task is to check if there is obvious speech that directly mentions or rephrases the values. If the story mentions phrases or sentences from values, that would be also counted as directly mentioned If it just reflects the value but does not rephrase the values, it is not considered as obvious speech. Here are some examples of obvious speech: [Example 1] Value: In your view, how often do the following things occur in this country’s elections: Journalists provide fair coverage of elections—Very often Speech: ”I think journalists do their best to cover things fairly” Explanation: The speech basically rephrases the value without providing supporting evidence [Example 2]: ……. You should output places where the values are directly mentioned. If the values are not directly spit out, do not output them!!! Here is the original story: story Here are the values to reflect in the story: values Now check if any parts of the stories directly mentioned the values. You can output reasoning or evidence and help you analyze. However, in the final answer, output speeches where the values are directly mentioned and the corresponding values. If there are no directly mentioned values, just output NO in the final answer. Follow the format: [Reasoning]: reasoning [Final answer]: [value]:[speech] [value]:[speech] …..
[…]
figureExplicit speech checkprompt
@@
Implicitness check prompt You will be provided a story and values reflected in the story. You will be also provided with some comments where the values are directly mentioned in the story. Each comment includes a value and a speech where the speech directly reflects the value. Your task is to refine the speech by rewriting the places where they directly mention the values. Instead of speaking out values directly, characters should reflect values implicitly by using anecdotes. Imagine the place is not a safe place to speak out, characters need to respect each other and not offend others. You need to make sure after rewriting, people are able to detect values by the speech. It cannot be too abstract. Here are some examples: [Example 1]: Value: How frequently do the following things occur in your neighborhood: Robberies?–Not frequently Original speech: “Well, at least we’re not dealing with robberies or anything in our area.” Refined speech: ”You know in other counties, cars got breaking in is common. At least we don’t have to worry about that.” [Example 2]: ….. Here is the original story: story Here are the values to reflect in the story: values Here are the comments that you need to refer to when refining the story: comments Please rewrite the speech one by one in the story. Do not miss any comments. For each rewritten speech, you need to follow this format: Value:[value] Original speech: [Speech in the comment] Refined speech: [You rewritten speech] Here is your output:
[…]
figureExplicit speech rewrite prompt
@@
Attitude detection prompt You will be given a scenario including background, conversation and actions in a location. You task is to detect the culture values reflected from the conversation. You will be provided a statement and you need to detect what attitude they hold to the value. For example, the statement is ”The people who don’t work turn lazy” and the options are ”disagree, agree, neither disagree nor agree”. And you need to read the story and choose one attitude from the options. The options could also describe something happening around them. For example, if the statement is ”How often does this happen in your neighborhood:Racist behaviors”, the options are ”Very often, not often, not at all often”, you need to choose one based on the story. Here is the scenario: story Here is the statement: statement Here are the options: options Now choose one option based on the story. Let’s think step by step! First, extract the speech that is related to the statement.Then analyse the attitude of those people from the speech. Finally, output the answer. Follow the format: [Related speech]: speeches [Analysis]: analysis of attitude [Answer]:answer
[…]
figureAttitude detection prompt. We show the zero-shot setting only. One-shot setting includes one demonstration and the no-reasoning setting does not include the step by step guidance.
@@
Value selection prompt You will be given a scenario including background, conversation and actions in a location. You task is to detect the culture values reflected from the conversation. You will be provided 15 options and you need to select number correct answers from the options. (Only number options are correct). You need to choose the whole option, for example: Being a housewife is just as fulfilling as working for pay–agree Make sure you include ’–’ in the answer. Here is the scenario: story Here are the options: options To think step by step: 1. You need to first detect what topics the story mentioned according to the options. You need to list the speech related to the topic. 2. Then for each topic, detect what values might be related to the topic. Sometimes, you might find multiple options that might be correct, for uncertain options, you should compare them. 3. Based on the previous detection, analyze what values are aligned with the story, you should always prioritize those values that are strongly related. 4. In the end, output the final answers only. Only choose the required number of values. Follow the format: [Topic]: topic1: speeches related to the topic ….. [Value detection]: topic1: what values are related to the topic [Reasoning]: Reasoning [Final answer]: [text] – [text] ….. ”””
[…]
figureValue selection prompt. We show the zero-shot setting only. One-shot setting includes one demonstration and the no-reasoning setting does not include the step by step guidance.
@@
Value selection prompt You will be provided with a story. Your goal is to summarize the main themes in the story. Please provide 4-5 sentences to summarize the ideas. You should be specifically looking for: topic themes. Please pay attention to the character’s responses and attitudes towards topic. Here is the story to summarize: story. You will be provided with a story. Your goal is to identify the most prevalent topic values in the story. For example, one religious value could be: I strongly believe in God and the afterlife. The value should be a complete sentence, it should not be a phrase like work-life balance. The value should hold a attitude or it reflects social phenomena, it should not be a overview of topic. For example, one social value could be: I think work is a duty towards the society. And one political value could be: I think voters are bribed in our election system. Please identify the values in this story: story by paying attention to how the characters in the story discuss topic. You need to generate 10 values in total, make sure values are specific and detailed. Only focus on one topic in one value. Do not mention several themes in one value. Here is the story: story Here is a story summary with some key ideas: summary. You could output reasoning before you output final answer. But in the end, your output should follow the format: [Final answer]: value1 value2 ….
[…]
figureValue extraction prompt.
@@
LLM evaluation prompt You will be given two sets of texts: a set of predicted values and a ground truth set of values. Your task is to determine how many of the ground truth values are fully represented in the predicted values. A ground truth value is considered correct if all of its components are meaningfully discussed in the predicted value, even if there is no exact 1-to-1 match. It could be many-to-1 match i.e. many values to 1 ground truth value. If the ground truth value is fully presented, score 1, if it is partially presented, score 0.5, if it is not mentioned at all, score 0. Additionally, provide a brief justification for your score, explaining which values were correctly or incorrectly represented, in the justification, you should explicitly mention which predicted values are related to the ground truth value. Here are the predictions: pred and ground truth: gt In the reasoning, if the ground truth value is fully represented, you need to point out which predicted value is related to it. Return the results in this format: [Reasoning]: [Ground truth 1]: [reasoning] ….. [Final answer]: [Ground truth 1]:1 [Ground truth 2]:0.5 ……
[…]
figureLLM evaluation prompt.

### v2

@@ Appendix D Prompts
In this section, we show all the prompts we use for story generation, experiments, and validation.
@@
Story generation prompt ##Identity You are an expert in culture understanding. ## Task You will be provided with 5 cultural values and a location where the conversation happens. Each of them follows the format [culture]–[value]. The first [culture] describes a statement or a situation, and [value] is how you agree with the culture or is the culture common or not. Your task is to generate a scene including conversations and actions among multiple people and the scene needs to reflect the culture values provided. Here are some requirements of the scene: 1. It cannot be too short. It should have multiple rounds of interaction among people. 2. It should not be too obvious. It cannot directly spit out or rephrase the values. 3. It cannot be easy to human to understand the culture values behind. 4. You do not need to follow the order of the values. You could mention the values multiple times through the conversation. Make sure the conversation flows well. 5. All the characters should follow the given values. There should not be contradictions between the character’s value and the given value. ##Input Here are the cultural values you should follow when generating: values Here is the pre-defined location of the scene: location ##Output Now using the cultural values to generate a story.
@@
Incorporation check prompt ##Identity You are an expert in culture understanding. ##Task You will be provided a story and values reflected in the story. Your task is to check if the story reflects values? The story does not mention the values directly. You will need some reasoning to analyze the story. ##Input Here is the original story: story Here are the values to reflect in the story: values ##Output For each value, output if the value is reflected and provide reasoning. In the end, output the values not reflected without the reasoning. Only output the exact and comprehensive value including "–" within it, do not rephrase! If all the values are reflected, just leave it blank. Follow the format: [Value]:[Reasoning, Yes/No] ….. Values not reflected: [Value]
@@
Missing values incorporation prompt ##Identity You are an expert in culture understanding. ##Task You will be provided a story and values which need to be reflected in the story. Your task is to refine the story to reflect the value provided. You cannot remove anything or replace existing speeches from the story, you can only add conversations to reflect the value. The refinement should flow with original story well. You cannot add new conversation randomly. ##Input Here is the original story: story Here are the values to reflect in the story: values ##Output Now refine the story.
@@
Consistency check prompt ##Identity You are an expert in culture understanding. ##Task You will be provided a story and values reflected in the story. Your task: for each value, check if all the characters agree with the value. If there is characters who does not agree with the value, you should output the character’s name and his speech, and why the speech does not align with the value. ##Input Here is the original story: story Here are the values to reflect in the story: values Now check the story and output if there is any contradiction. You can output reasoning to help you analyze. However, in the end, only output where is the contradiction one by one. If there isn’t a contradiction, just reply NO. Otherwise, reply where is the contradiction. ##Output Follow the format strictly, do not change the format, output exact values from the values provided, and do not rephrase: [Reasoning]: [Value–attitude]:[If all the speeches are aligned with the value] [Contradictions]: [Value–attitude]: [*character name*:speech] ……
@@
Consistency resolve prompt ##Identity You are an expert in culture understanding. ##Task You will be provided a story and values which need to be reflected in the story. However, the story includes some contradiction where characters do not agree on certain values. You will be provided where is the contradiction. The contradiction includes 3 parts: 1. Correct value to follow 2. Character name 3. Character speech Your task is to replace the speech mentioned in the contradictions with a new speech to make sure the speech is aligned with the values The refinement should flow with the original story well. You cannot add new conversations randomly. ##Input Here is the original story: story Here are the contradictions: contradiction Ignore the original character’s speech. Directly write a new speech that reflects the value. ##Output Here is the rewritten speech:
@@
Implicitness check prompt ##Identity You are an expert in culture understanding. ##Task You will be provided a story and values reflected in the story. Your task is to check if there is obvious speech that directly mentions or rephrases the values. If the story mentions phrases or sentences from values, that would be also counted as directly mentioned If it just reflects the value but does not rephrase the values, it is not considered as obvious speech. Here are some examples of obvious speech: [Example 1] Value: In your view, how often do the following things occur in this country’s elections: Journalists provide fair coverage of elections—Very often Speech: "I think journalists do their best to cover things fairly" Explanation: The speech basically rephrases the value without providing supporting evidence [Example 2]: ……. You should output places where the values are directly mentioned. If the values are not directly spit out, do not output them!!! ##Input Here is the original story: story Here are the values to reflect in the story: values ##Output Now check if any parts of the stories directly mentioned the values. You can output reasoning or evidence and help you analyze. However, in the final answer, output speeches where the values are directly mentioned and the corresponding values. If there are no directly mentioned values, just output NO in the final answer. Follow the format: [Reasoning]: reasoning [Final answer]: [value]:[speech] [value]:[speech] …..
@@
Implicitness rewrite prompt ##Identity You are an expert in culture understanding. ##Task You will be provided a story and values reflected in the story. You will be also provided with some comments where the values are directly mentioned in the story. Each comment includes a value and a speech where the speech directly reflects the value. Your task is to refine the speech by rewriting the places where they directly mention the values. Instead of speaking out values directly, characters should reflect values implicitly by using anecdotes. Imagine the place is not a safe place to speak out, characters need to respect each other and not offend others. You need to make sure after rewriting, people are able to detect values by the speech. It cannot be too abstract. Here are some examples: [Example 1]: Value: How frequently do the following things occur in your neighborhood: Robberies?–Not frequently Original speech: “Well, at least we’re not dealing with robberies or anything in our area.” Refined speech: "You know in other counties, cars got breaking in is common. At least we don’t have to worry about that." [Example 2]: ….. ##Input Here is the original story: story Here are the values to reflect in the story: values Here are the comments that you need to refer to when refining the story: comments ##Output Please rewrite the speech one by one in the story. Do not miss any comments. For each rewritten speech, you need to follow this format: Value:[value] Original speech: [Speech in the comment] Refined speech: [You rewritten speech] Here is your output:
@@
Attitude detection prompt ##Identity You are an expert in culture understanding. ##Task You will be given a scenario including background, conversation and actions in a location. You task is to detect the culture values reflected from the conversation. You will be provided a statement and you need to detect what attitude they hold to the value. For example, the statement is "The people who don’t work turn lazy" and the options are "disagree, agree, neither disagree nor agree". And you need to read the story and choose one attitude from the options. The options could also describe something happening around them. For example, if the statement is "How often does this happen in your neighborhood:Racist behaviors", the options are "Very often, not often, not at all often", you need to choose one based on the story. ##Input Here is the scenario: story Here is the statement: statement Here are the options: options ##Output Now choose one option based on the story. Let’s think step by step! First, extract the speech that is related to the statement.Then analyse the attitude of those people from the speech. Finally, output the answer. Follow the format: [Related speech]: speeches [Analysis]: analysis of attitude [Answer]:answer
@@
Value selection prompt ##Identity You are an expert in culture understanding. ##Task You will be given a scenario including background, conversation and actions in a location. You task is to detect the culture values reflected from the conversation. You will be provided 15 options and you need to select number correct answers from the options. (Only number options are correct). You need to choose the whole option, for example: Being a housewife is just as fulfilling as working for pay–agree Make sure you include ’–’ in the answer. ##Input Here is the scenario: story Here are the options: options ## Instruction To think step by step: 1. You need to first detect what topics the story mentioned according to the options. You need to list the speech related to the topic. 2. Then for each topic, detect what values might be related to the topic. Sometimes, you might find multiple options that might be correct, for uncertain options, you should compare them. 3. Based on the previous detection, analyze what values are aligned with the story, you should always prioritize those values that are strongly related. 4. In the end, output the final answers only. Only choose the required number of values. ##Output Follow the format: [Topic]: topic1: speeches related to the topic ….. [Value detection]: topic1: what values are related to the topic [Reasoning]: Reasoning [Final answer]: [text] – [text] ….. """
@@
Value extraction prompt ##Identity You are an expert in culture understanding. ##Task You will be provided with a story. Your goal is to identify the most prevalent topic values in the story. For example, one religious value could be: I strongly believe in God and the afterlife. The value should be a complete sentence, it should not be a phrase like work-life balance. The value should hold a attitude or it reflects social phenomena, it should not be a overview of topic. For example, one social value could be: I think work is a duty towards the society. And one political value could be: I think voters are bribed in our election system. Please identify the values in this story: story by paying attention to how the characters in the story discuss topic. You need to generate 10 values in total, make sure values are specific and detailed. Only focus on one topic in one value. Do not mention several themes in one value. ##Input Here is the story: story Here is a story summary with some key ideas: summary. ##Output You could output reasoning before you output final answer. But in the end, your output should follow the format: [Final answer]: value1 value2 ….
@@
LLM evaluation prompt ##Identity You are an expert in culture understanding. ##Task You will be given two sets of texts: a set of predicted values and a ground truth set of values. Your task is to determine how many of the ground truth values are fully represented in the predicted values. A ground truth value is considered correct if all of its components are meaningfully discussed in the predicted value, even if there is no exact 1-to-1 match. It could be many-to-1 match i.e. many values to 1 ground truth value. If the ground truth value is fully presented, score 1, if it is partially presented, score 0.5, if it is not mentioned at all, score 0. Additionally, provide a brief justification for your score, explaining which values were correctly or incorrectly represented, in the justification, you should explicitly mention which predicted values are related to the ground truth value. ##Input Here are the predictions: pred and ground truth: gt In the reasoning, if the ground truth value is fully represented, you need to point out which predicted value is related to it. ##Output Return the results in this format: [Reasoning]: [Ground truth 1]: [reasoning] ….. [Final answer]: [Ground truth 1]:1 [Ground truth 2]:0.5 ……